OpenStudy (anonymous):

Solve the equation below. Please show all of your work. log2(1/64)=1/3-x Solve the equation below. Please show all of your work. log2(1/64)=1/3-x @Mathematics

OpenStudy (anonymous):

x= -log2(1/64) + 1/3 x= 19/3

OpenStudy (anonymous):

how did you get this?

jhonyy9 (jhonyy9):

because loga +logb=log(a*b) so this is property of logaritms

jhonyy9 (jhonyy9):

but i think that this not is correct because so 1/64 = 1/2(6) like 2 on exponent 6 so log2(

jhonyy9 (jhonyy9):

so will be log2(2) on exponent minus 6

jhonyy9 (jhonyy9):

than will be -6log2(2) = -6

jhonyy9 (jhonyy9):

than -6=1/3 -x so -18=1-3x -19=-3x 19=3x x=19/3

OpenStudy (anonymous):

it's all numbers... log2(1/64) is a number (-6) just like 1/3 is a number (.3333) think of it as -6=.333-x and solve for x

OpenStudy (anonymous):

log2(1/64) = log(1/64)/log(2)